OpenStudy (anonymous):

Really stuck on this one here! What is the approximate area of the output piston if the diameter is 1.4 m? Use the formula: A = π r2.

jagr2713 (jagr2713):

well when its a diameter that means you divide it by 2 so if the diameter was 4 then u would divide it by 2 and get 2, you get it @42cookiethere

OpenStudy (anonymous):

radius is diameter/2 work from that :)

OpenStudy (chaise):

The diametre is usually 2 x the radius. In this case, your radius is 0.7m

OpenStudy (alexandervonhumboldt2):

use the formula: pi=3.14. r=1/2*1.4=0.7 substitue: A=0.7*0.7*3.14=?

OpenStudy (anonymous):

okay ummm...I have no idea, haha! sorry this stuff is mind boggling for me!

jagr2713 (jagr2713):

so if the diameter is 1.4 then u divide by 2 and get 0.7 then b.c its says r^@ you multiply it by itself and get 0.49 so R=? what @42cookiethere

OpenStudy (anonymous):

2 by 2 almost like algbra x variable

jagr2713 (jagr2713):

you get it @42cookiethere

jagr2713 (jagr2713):

if you plug it in the formula like this \[A=\pi 1.4 ^{2}\] so if you plug it in u get

OpenStudy (anonymous):

I don't get it!!!!! ahhh!

jagr2713 (jagr2713):

so you divide 1.4/2 what do u get @42cookiethere

OpenStudy (anonymous):

OMG 0.7! there! Thank you!

jagr2713 (jagr2713):

if it was radius then u dont divide, so you plug it in A=pi 0.7^2 so u multiply 0.7 *0.7 and then u multiply it with 3.14, what do u get

OpenStudy (anonymous):

1.5386

jagr2713 (jagr2713):

Correct

jagr2713 (jagr2713):

see you did it :D

OpenStudy (anonymous):

Phew! thankyou! sorry, I just got impatient! haha! thanks!

jagr2713 (jagr2713):

anytime